UNDERSTANDING THE MOLECULAR PROPERTIES OF ClF_{n} (n =1-7) SPECIES: AN APPLICATION OF THE RECOUPLED PAIR BONDING MODEL FOR HYPERVALENT BONDS

\underline{L. CHEN}, D. E. WOON, and T. H. DUNNING, Jr., Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

Recently, new insight into the nature of hypervalent behavior led us to develop a model called recoupled pair bonding. In this model, two hypervalent bonds can be formed by decoupling a valence p^{2} or s^{2} electron pair. However, energy must be expended to decouple an electron pair, and the first bond is weakened as a consequence. The recoupled pair bonding model has been proven successful in our initial study of the SF_{n} (n =1-7) species. To further examine the applicability of this new model, this study explored the molecular properties of the ClF_{n} (n =1-7) series. Optimized ground state structures, bond energies, and spectral properties of these molecules were obtained by employing high level ab initio calculations [MRCI, CCSD(T)] with correlation consistent basis sets. Because of recoupled pair bonding, there are unanticipated low-lying excited states such as ClF (^{3}Π) and ClF_{2} (^{2}Π,^{4}Σ). We also systematically explored the bond formation processes, adding F atoms one at a time to the optimized ClF_{n} (1 ≤ n ≤ 6 ) molecules. We find the bond energies for F addition to form ClF_{2}, ClF_{4}, and ClF_{6} are much lower than those leading to ClF, ClF_{3} and ClF_{5}. This oscillating trend is analogous to what is seen in the SF_{n} species, though the bond energies of the SF_{n} species are considerably greater than the ones for ClF_{n}. The lower bond energies of the even n species in the ClF_{n} series reflects the cost of decoupling paired electrons of the central atom, and the difference between ClF_{n} and SF_{n} reflects the fact that more energy is needed to decouple each of the 3p^{2} pairs of electrons of Cl than the single 3p^{2} pair of S. This behavior and other trends observed in ClF_{n} species demonstrate the improved predictive ability of the recoupled pair bonding model over other models for describing hypervalent bonding.